40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 14 19:41:46 2017, Jamie, Update, CDS, front-end/DAQ network down for kernel upgrade, and timing errors 
    Reply  Wed Aug 16 17:05:53 2017, Jamie, Update, CDS, front-end/DAQ network down for kernel upgrade, and timing errors 2017-08-16-163725_1366x495_scrot.png
       Reply  Wed Aug 16 17:14:02 2017, Koji, Update, CDS, front-end/DAQ network down for kernel upgrade, and timing errors 
          Reply  Wed Aug 16 18:01:28 2017, Jamie, Update, CDS, front-end/DAQ network down for kernel upgrade, and timing errors 
             Reply  Wed Aug 16 18:06:01 2017, Koji, Update, CDS, front-end/DAQ network down for kernel upgrade, and timing errors 
       Reply  Wed Aug 16 18:50:58 2017, Jamie, Update, CDS, front-end/DAQ network down for kernel upgrade, and timing errors 2017-08-16-184910_1394x488_scrot.png
          Reply  Mon Aug 28 16:20:00 2017, gautam, Update, CDS, 40m files backup situation 
             Reply  Mon Aug 28 17:13:57 2017, ericq, Update, CDS, 40m files backup situation 
             Reply  Fri Sep 15 15:54:28 2017, gautam, Update, CDS, FB wiper script 
                Reply  Mon Sep 18 17:17:49 2017, gautam, Update, CDS, FB wiper script 
                   Reply  Mon Sep 18 17:30:54 2017, Chris, Update, CDS, FB wiper script wiper.pl
                      Reply  Mon Sep 18 17:51:26 2017, gautam, Update, CDS, FB wiper script perlDiff.png
                         Reply  Mon Sep 18 18:40:34 2017, gautam, Update, CDS, FB wiper script 
             Reply  Tue Sep 26 15:55:20 2017, gautam, Update, CDS, 40m files backup situation 
                Reply  Thu Sep 28 10:33:46 2017, gautam, Update, CDS, 40m files backup situation 
                   Reply  Thu Sep 28 11:13:32 2017, jamie, Update, CDS, 40m files backup situation 
                      Reply  Thu Sep 28 23:47:38 2017, gautam, Update, CDS, 40m files backup situation 
                         Reply  Fri Sep 29 11:07:16 2017, gautam, Update, CDS, 40m files backup situation 
                            Reply  Thu Oct 5 13:58:26 2017, gautam, Update, CDS, 40m files backup situation 
                               Reply  Fri Oct 6 12:46:17 2017, gautam, Update, CDS, 40m files backup situation 
                                  Reply  Sat Oct 28 00:36:26 2017, gautam, Update, CDS, 40m files backup situation - ddrescue 415E2F09-3962-432C-B901-DBCB5CE1F6B6.jpegBFF8F8B5-1836-4188-BDF1-DDC0F5B45B41.jpeg
Message ID: 13215     Entry time: Wed Aug 16 17:05:53 2017     In reply to: 13205     Reply to this: 13216   13219
Author: Jamie 
Type: Update 
Category: CDS 
Subject: front-end/DAQ network down for kernel upgrade, and timing errors 

The CDS system has now been up moved to a supposedly more stable real-time-patched linux kernel (3.2.88-csp) and RCG r4447 (roughly the head of trunk, intended to be release 3.4).  With one major and one minor exception, everything seems to be working:

The remaining issues are:

  • RFM network down.  The IOP models on all hosts on the RFM network are not detecting their RFM cards.  Keith Thorne thinks that this is because of changes in trunk to support the new long-range PCIe that will be used at the sites, and that we just need to add a new parameter to the cdsParameters block in models that use RFM.  Him and Rolf are looking into for us.
  • The 3.2.88-csp kernel is still not totally stable.  On most hosts (c1sus, c1ioo, c1iscex) it seems totally fine and we're able to load/unload models without issue.  c1iscey is definitely problematic, frequently freezing on module unload.  There must be a hardware/bios issue involved here.  c1lsc has also shown some problems.  A better kernel is supposedly in the works.
  • NDS clients other than DTT are still unable to raise test points.  This appears to be an issue with the daqd_rcv component (i.e. NDS server) not properly resolving the front ends in the GDS network.  Still looking into this with Keith, Rolf, and Jonathan.

Issues that have been fixed:

  • "EDCU" channels, i.e. non-front-end EPICS channels, are now being acquired properly by the DAQ.  The front-ends now send all slow channels to the daq over the MX network stream.  This means that front end channels should no longer be specified in the EDCU ini file.  There were a couple in there that I removed, and that seemed to fix that issue.
  • Data should now be recorded in all formats: full frames, as well as second, minute, and raw_minute trends
  • All FE and DAQD diagnostics are green (other than the ones indicating the problems with the RFM network).  This was fixed by getting the front ends models, mx_stream processes, and daqd processes all compiled against the same version of the advLigoRTS, and adding the appropriate command line parameters to the mx_stream processes.
ELOG V3.1.3-